OpenStudy (anonymous):

Find the solution to each system by the addition (elimination) method. Check your answer for exercises 23-26. 5s+9r=6 4s-3t=15

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

I'm Still suck on how to figure out how to get the -1

OpenStudy (uri):

@Luigi0210 I'm confused why is it r and t.

OpenStudy (luigi0210):

Uri. different variables can be used

OpenStudy (uri):

Well explain her i also wanna learn..

OpenStudy (anonymous):

its suppsoe to be 5s+9t=6 4s-9t=15

OpenStudy (luigi0210):

There Uri, you can handle ir now

OpenStudy (luigi0210):

*it

OpenStudy (uri):

Told yea..:p

OpenStudy (uri):

Okay so just simply add it 5s+9t=6 4s-9t=15 ======== Remember +9t -9t will be cancelled. @klovelace

OpenStudy (anonymous):

Ive got that part down I just dont understand where to get the -1

OpenStudy (uri):

-1 hue? just do what i'm telling..and you'll know honey

OpenStudy (uri):

So what's left after adding them? 5s+9t=6 4s-9t=15 ========